trail mix has 10 grams of raisins and 6 grams of sunflower seeds, how many sunflower seeds is in 40 grams of trail mix? Need to make a ratio chart as well as a tape chart.
I have tried for 3 hours but I can't get it to work. I tried doing half my chart
5 10 15 20 25 30 35 raisins
3 6 9 12 15 18 21 sunflower seeds
8 16 24 32 40 48 56
to me the answer should be 15. but I don't know if this is right and then how to get it in a tape form

Answer by rothauserc(4718)

10x +6x = 40
5x +3x = 20
8x = 20
x = 5/2 = 2.5
there are 6*(5/2) grams of sunflower seeds, so
there are 15 grams of sunflower seeds
does this help you do the tape?